OpenStudy (anonymous):

HELP!!!!!!!! FAN AND MEDAL!!!!!!! The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (anonymous):

@InYourHead

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

This is a first degree equation, where all the variables are to the power of one (A^1, etc.) You know one of the variables, which is A. You need to know the second variable (R) to answer the question: "how many pounds of Robusta beans should be put into the mixture?". When you substitute the value A (1000) into the equation, it looks like this: \[700\times(1000) + 1200R = 1,000,000\] Now you can solve for R: \[700000 + 1200R = 1,000,000\] \[1200R=300000\] \[R=250\] Therefore the amount needed to put in the mixture of Robusta beans is 250 pounds. For the second part, just explain in your own words my first paragraph, then explain each step to solve for R. Hope I helped out :)

OpenStudy (anonymous):

thank you